OpenStudy (anonymous):

* Financial Algebra * [Work Is Shown] Zoe purchased an entertainment center for $3,257 using a 12-month deferred payment plan with an interest rate of 28.05%. She did not make any payments during the deferment period. What will the total cost of the entertainment center be if she must pay it off within four years after the deferment period? $3,257.00 $4,297.69 $7,195.68 $8,994.60

OpenStudy (anonymous):

I plug this in B=3257(1+(0.023375))^(12*1.5)

OpenStudy (anonymous):

I got 4936.79

OpenStudy (amistre64):

12 months defered is not 1.5 years

OpenStudy (amistre64):

i can use my formula, adjusting for the defered period as:\[B_n=B_ok^{d+n}-P\frac{1-k^n}{1-k}\] d = 12, n = 4 years of 12 months each = 48 solving for P*48 tells how much was paid in total

OpenStudy (amistre64):

P = 3257k^(12+48)(1-k)/(1-k^48) 48 P = 48*3257k^(12+48)(1-k)/(1-k^48), k=1+.2802/12 http://www.wolframalpha.com/input/?i=48*3257k%5E%2812%2B48%29%281-k%29%2F%281-k%5E48%29%2C+k%3D1%2B.2802%2F12

OpenStudy (amistre64):

well, .2805 but thats in the ball park nonetheless

OpenStudy (anonymous):

so the answer is 7,195.68

OpenStudy (anonymous):

you 12/4=48 ?

OpenStudy (anonymous):

wait that was a stupid queation

OpenStudy (amistre64):

12 months of (d)eferred 4 years of 12 payments a year is 48 payments total

OpenStudy (anonymous):

oh you multiplied . Do I add it in my formula ?

OpenStudy (amistre64):

the formula you presented just gives you a defered balance after a year and a half, ... its rather incomplete as far as a solution goes

OpenStudy (amistre64):

the key is to determine the payment amount: 149.91 then multiply it by 48 payments to see how much you piad ....

OpenStudy (anonymous):

would 7,195.68 be my answer .

OpenStudy (amistre64):

of course

OpenStudy (anonymous):

ok , Thank you alot again lol

OpenStudy (amistre64):

yw